A sample of sewage from a town is found to have a BOD5 of 210 mg/L and a k value of 0.11/day. Estimate the ultimate biological demand for this sewage (mg/L).
Answer #1

ANSWER :-

BOD (at a particular time) 5 = L​​​​​​0 ( 1 - e-kt​​​​​​ )

where,

1. k is the speed of BOD reaction

2. L0 is the ultimate biological demand

3. ' t ' is the time period of 5 days

BOD5 = L0 × [ 1 - e(-0.11)×5 ]

210 mg/L = L0 × ( 1 - 0.5769 )

210 = L0 × (0.4231)

So, L0 = 210 / 0.4231 = 496.33 mg/L

CONCLUSION :- The ultimate Biological Oxygen Demand for the sewage is 496.33 mg/L.